OpenStudy (anonymous):

What is the domain and range of: 1) sqrt of x^2 -4 2) sqrt of 4-x2

OpenStudy (anonymous):

\[x^2 - 4 \ge 0 \] \[x\ge 2\]

OpenStudy (anonymous):

\[4 -x^2 \ge 0 \] \[ 2\ge x\]

OpenStudy (anonymous):

how about teh ranges

OpenStudy (anonymous):

how come for 4-x^2, teh domain isn't all real numbers?

OpenStudy (mimi_x3):

because its in a square root xD

OpenStudy (anonymous):

range y^2 = x^2 - 4 (sqrt(y^2 + 4)) = x all real numbers y^2 = 4 -x^2 x = sqrt(4 -y^2) \[4 - y^2 \ge 0 \] \[ 2\ge y \]

OpenStudy (anonymous):

thanks Mim_x3 ah I feel so honored she replied to one of my questions ..

OpenStudy (mimi_x3):

LOL, sorry for interrupting in you answering the question Mimi*

OpenStudy (anonymous):

No it's an honor

OpenStudy (mimi_x3):

LOL

OpenStudy (anonymous):

i know this sounds dumb but if it write taht \[y \ge 2\] is taht teh same thing as what u wrote: \[2 \ge y\]

OpenStudy (anonymous):

i mean y is less tahn 2

OpenStudy (anonymous):

\[y \le 2 \]

OpenStudy (anonymous):

no it isn't first is y is always greater than 2 second is y is less than 2

OpenStudy (mimi_x3):

lols , no you only write y is less then 2 when its a minus xD

OpenStudy (anonymous):

Thank you both! =)

OpenStudy (mimi_x3):

yw (: